OpenStudy (anonymous):

Through how many radians does the minute hand of a clock turn in 10 minutes?

OpenStudy (unklerhaukus):

\[\frac{10min}{60min}=\frac{r}{2\pi}\]

OpenStudy (anonymous):

why is it radius over 2pie?

OpenStudy (unklerhaukus):

\[r=2\pi\frac{10min}{60min}=\frac\pi3\]

OpenStudy (anonymous):

ooh nevermind

OpenStudy (anonymous):

oh

OpenStudy (unklerhaukus):

because there are 2π radiens in a complete revolution

OpenStudy (anonymous):

^

OpenStudy (anonymous):

thanks

OpenStudy (unklerhaukus):

2π radiens = 360°